# PATCH - Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Mathematics Proficiency

### v1

@@ 3. Answer:
You are given a math question written in LaTeX format.
@@
You are given a math question written in LaTeX format. Instructions: 1. Type of question: Is it multiple choice, free text response, or drawing? 2. Think step by step, and describe your thought process and reasoning. 3. Answer: - For multiple choice: [selected answer key]. - For free-text response: [provide your short answer]. - For drawing: [describe clearly the steps to complete the drawing]. - If uncertain, make an educated guess.
@@
You are given a math question written in LaTeX format.
@@ Appendix B Example Responses
2. First, I notice that the number line is labeled from 26 to 28. This means that the value of $K$ must be between 26 and 28.

### v2

@@ 3. Answer:
You are given a mathematics question written in LaTeX format.
@@
You are given a mathematics question written in LaTeX format. Instructions: 1. Type of question: Is it multiple choice, free text response, or drawing? 2. Think step by step, and describe your thought process and reasoning. 3. Answer: - For multiple choice: [selected answer key]. - For free-text response: [provide your short answer]. - For drawing: [describe clearly the steps to complete the drawing]. - If uncertain, make an educated guess.
@@
You are given a mathematics question written in LaTeX format.
@@ Appendix B Example Responses
2. First, I notice that the number line is labelled from 26 to 28. This means that the value of $K$ must be between 26 and 28.